An ice cream scoop scoops out ice cream spheres with radius 1 inch. How many scoops are needed to fill an ice cream cone with radius 2 inches and height 5 inches?
Answer by Cromlix(4381)

Hi there,
Volume of a cone = 1/3 x pi x radius^2 x height
Volume = 1/3 x pi x 2^2 x 5
Volume = 20.94 inches^3 (2 decimal places)
Volume of a sphere (scoop) 4/3 x pi x radius^3
Volume = 4/3 x pi x 1^3
Volume = 4.19 inches^3 (2 decimal places)
20.94/4.19
= 4.99. (2 decimal places)
or about 5 scoops.
Hope this helps :-)